Isabel, a calendar-year taxpayer, uses the cash method of accounting for her sole proprietorship. In late December she receivc. Based on requirements a and b, should Isabel pay the $41,000 bill in December or January? O December January
Answer #1

a) After Tax Cost, if Isabel pays the bill $41000 in December :

Bill Amount = $ 41000

Less: Tax Saving = 41000*37% = $15170

After Tax Cost = $25830

Note: On payment of an expense, there will be saving in tax which will reduce the effective cost of that expense.

b) After Tax Cost, if Isabel pays the bill $41000 in January :

Bill Amount = $ 41000

Less: Tax Saving = 41000*37% = $15170

Less: Return on Investment @ 6% for one month  = 41000*6%*1/12 = $205

After Tax Cost = $25625

Note1 : On payment of an expense, there will be saving in tax which will reduce the effective cost of that expense.

Note 2 : We are assuming that Isabel will pay the bill at the end of January and therefore he will be eligible for Interest on Investment for one month .

Note 3: Return on Investment is an income and will reduce the net Cost of the expense.

Note 4: we are given the rate of Return of investment as 6% so we are assuming that is for annum, therfore, we divide it by 12 to get one month rate of return . Alternatively we can assume it for one month then the Return of investment anount would be 41000*6% = $ 2460 ( Although this alternative is not practically possible)

Note 5. We are given 6% rate of return after considering tax effect. Therefore we will deduct it from after tax effective cost of bill.

C) On the based of above wo alternative a) and b) , it is advisable for Isabel to choose Option b) that is pay in January , it will cost less to her by $205.